. A battery that has q=10000 C

q=10000 C and ΔV=12 V

ΔV=12 V. Find the total electric potential energy.


1
Expert's answer
2022-03-28T14:28:31-0400

Answer

the total electric potential energy

"U=\frac{qΔV}{2}\\=\frac{10000*12}{2}\\=60000J"